A rectangle is four times as long as it is wide .if the perimeter is 50 inches find the length.
Answer by ewatrrr(24785)

 
Hi,
.
L + w = 25in
4w + w = 25
5w = 25
w = 5" and L = 20"
And...checking
L + w = 25in